OpenStudy (anonymous):

Find cot θ if csc θ = sqrt17/4 and tan θ > 0.

OpenStudy (anonymous):

csc θ=1/sin θ= sqrt17/4 sin θ=4/sqrt17 then you should use the formula sin^2 θ+cos^2 θ=1. can you do that??

OpenStudy (anonymous):

um i think so! So, 4/sqrt17^(2) + cos^(2)θ=1... then would I use the sin^2θ + cos^2θ = 1 formula?

OpenStudy (anonymous):

you must use the\[\sin ^{2}θ+\cos ^{2}θ=1\] ok??

OpenStudy (anonymous):

yeah! so (4/sqrt17)^2 + (4/sqrt17)^2-1.... would that be right?

OpenStudy (anonymous):

(=1)

OpenStudy (anonymous):

\[\ (\frac{ 4 }{ \sqrt{17} })^{2}+\cos ^{2}θ=1\] \[\frac{ 16 }{ 17 }+\cos ^{2}θ=1\] cos^2 θ=1/17 and cos θ=1/sqrt17 can you find cot θ??

OpenStudy (anonymous):

cot θ = (1/sqrt17) / (4/sqrt17) = 1/4?

OpenStudy (anonymous):

ok. you're correct :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

welcome :)